OpenStudy (anonymous):

Using (f circle g)(x). a) f(x) = x + 5 g(x) = x - 8 b) f(x) = 2x - 3 g(x) = x + 4 c) f(x) = 3x - 1 g(x) = 9x^2 - 1 d) f(x) = 2x^2 - x -1 g(x) = 4x^2 -1 e) f(x) = x^3 - 1 g(x) = x^2 - 1

ganeshie8 (ganeshie8):

use the definition : \(\large (f\circ g)(x) = f\left(g(x)\right)\)

ganeshie8 (ganeshie8):

a) \(f(x) = x + 5\) \(\color{red}{g(x) = x - 8}\) \(\large (f\circ g)(x) = f\left(\color{Red}{g(x)}\right)\) \(\large ~~~~~~~~~~~~~ = f\left(\color{red}{x-8}\right)\) \(\large ~~~~~~~~~~~~~ = ?\)

ganeshie8 (ganeshie8):

just replace `x` by `x-8` in f(x)

OpenStudy (anonymous):

ok thank you! :D

ganeshie8 (ganeshie8):

what did u get after that step ?